OpenStudy (anonymous):

Estimate the length of the radius, if the area of a circle is 275 mm2. A. 9.78 mm B. 24.97 mm C. 9.36 mm D. 7.87 mm

OpenStudy (ipwnbunnies):

Area = pi * r^2 Where 'r' is the radius. r = \[\sqrt{area/\pi}\]

OpenStudy (anonymous):

A.